find the solutions of equation


(x-1/x)1/2 + (1-1/x)1/2 = x

Grade:11

2 Answers

Shashwat Tiwari
12 years ago
(x-1/x)1/2 = x-(1-1/x)1/2
squaring both the sides..........
x-1/x = x2 + 1 - 1/x - 2(x2-x)1/2 ...................
x2-x+1=2(x2-x)1/2.....................
now, put x2-x=t.......
t+1=2(t)1/2
again square both sides.............
(t-1)2=0
t=1
Hence, x2-x=1
x=(1±(5)1/2)/2..............Cool
jagdish singh singh
12 years ago